OpenStudy (anonymous):

translate sentence into an equation three less than one-half a number is-71

OpenStudy (anonymous):

or 3 - 0.5 x = -71

OpenStudy (anonymous):

i dont get it how u get 1/5

OpenStudy (anonymous):

one half.... another words 1/5.

OpenStudy (anonymous):

oh but these are my answer choices -1/2n+2=-71 2n-1/2=-71 1/2n-3=-71

OpenStudy (anonymous):

OMG. I'm so sorry, I just noticed... It would be the last one 1/2n - 3 = -71.

OpenStudy (anonymous):

ok thanks lots